OpenStudy (thatawayz):

Have a Geometry: Basic Trig Question, Not seeking to get answers from you, rather a step by step guide

OpenStudy (thatawayz):

Find the length of the base of the following pyramid, given the height of the pyramid is 211 meters and the angle of elevation of the base of the pyramid is 62°. Round to the nearest whole number.

OpenStudy (thatawayz):

How do i set this up

OpenStudy (anonymous):

So it is 90 pyramid right?

OpenStudy (anonymous):

*90 degree

OpenStudy (thatawayz):

Hi again lol, um yes it is a 90 degree pyramid

OpenStudy (anonymous):

So we have 

OpenStudy (thatawayz):

whoa, that was quick, yes that is the one

OpenStudy (anonymous):

Well we are not working with the hypotenuse, so which trig function features opposite and adjacent?

OpenStudy (thatawayz):

we are working with the tangent function right?

OpenStudy (anonymous):

yes, we are

OpenStudy (thatawayz):

sorry glitchy

OpenStudy (thatawayz):

so \[\tan 52 Degrees = 211/a\]

OpenStudy (thatawayz):

and whatever the adjacent is, the base of the pyramid is double

OpenStudy (anonymous):

Well, close \[\large \sf Tan62=\frac{211}{Base}\] So do the trick and change the a and c

OpenStudy (thatawayz):

oh, im sorry about that

OpenStudy (thatawayz):

Base= 211/tan 62

OpenStudy (anonymous):

Correct. So now we just crunch the numbers

OpenStudy (thatawayz):

i wrote it down in correct format, typing on phone is a bit weird

OpenStudy (thatawayz):

alrighty, so how would i "crunch the numbers"

OpenStudy (anonymous):

Well, I meant calculate it. You know do \(\huge \sf \frac{211}{tan62}\)

OpenStudy (thatawayz):

ahhh, ok

OpenStudy (thatawayz):

tan of 62 degrees is roughly 1.88, on phone calculator

OpenStudy (anonymous):

Ye that looks good to me

OpenStudy (thatawayz):

so 211+1.88

OpenStudy (thatawayz):

not plus, divided

OpenStudy (anonymous):

Yes 211/1.88

OpenStudy (thatawayz):

112.234

OpenStudy (anonymous):

That would be your answer

OpenStudy (thatawayz):

so double that is 224, the answer is c

OpenStudy (anonymous):

Why are we doubling it again?

OpenStudy (thatawayz):

seriously, thankyou for helping me with these, it means alot to me, im not corrently in the best place, graduation status and this really helps

OpenStudy (thatawayz):

the triangles 90 degree angle is in the center of the pyramid, finding the tan gives me half of the whole base of the pyramid

OpenStudy (thatawayz):



OpenStudy (anonymous):

Ahh, I see. That looks right, then

OpenStudy (anonymous):

And no problem, I'm happy to help.

OpenStudy (thatawayz):

Should i keep question open or close it?

OpenStudy (anonymous):

You have had your question resolved, so it would be time to close it.